The combined blood supply for an average man and woman is 9liters. The average man has 2.4liters more blood than the average woman. How much blood does an average man have? ( answer is 5.7liters - show your solution )
Answer by macston(5194)

Average woman blood supply =W; average man blood supply=M=W+2.4l
combined: M+W=9l Substitute for M
W+2.4l+W=9l
2W+2.4l=9l Subtract 2.4l from each side
2W=9l-2.4l
2W=6.6l Divide each side by 2
W=3.3l
M=W+2.4l
M=3.3l+2.4l=5.7l ANSWER Average man has 5.7 liters of blood.
CHECK
Average man's supply + average woman's supply= 9 liters
5.7 liter + 3.3 liter = 9 liter
9 liter= 9 liter